# Topological phases in discrete stochastic systems

### v1

@@ Abstract
Topological invariants have proved useful for analyzing emergent function as they characterize a property of the entire system, and are insensitive to local details, disorder, and noise. They support edge states, which reduce the system response to a lower dimensional space and offer a mechanism for the emergence of global cycles within a large phase space. Topological invariants have been heavily studied in quantum electronic systems and been observed in other classical platforms such as mechanical lattices. However, this framework largely describes equilibrium systems within an ordered crystalline lattice, whereas biological systems are often strongly non-equilibrium with stochastic components. We review recent developments in topological states in discrete stochastic models in 1d and 2d systems, and initial progress in identifying testable signature of topological states in molecular systems and ecology. These models further provide simple principles for targeted dynamics in synthetic systems or in the engineering of reconfigurable materials. Lastly, we describe novel theoretical properties of these systems such as the necessity for non-Hermiticity in permitting edge states, as well as new analytical tools to reveal these properties. The emerging developments shed light on fundamental principles for non-equilibrium systems and topological protection enabling robust biological function.
@@ I Introduction
Understanding how the underlying components of proteins, RNAs, metabolites, and other biologically relevant molecules give rise to biological function would enable precise targeting of interventions to enhance health. However, we still do not have a good theory for how structure leads to function in these complex and noisy systems. As just one example, despite successful sequencing of the human genome, it remains challenging to predict the behavior of resulting proteins and macromolecules due to the large space of possible configurations and reactions causing transitions between them [1, 2]. This large phase space of possible configurations also renders unfeasible exhaustive searches using other approaches like experiment or numerical simulation [3], underscoring the need for simple conceptual methods to provide insight [4, 5, 6].
Drawing on recent theoretical advances in topology, molecular biology, and non-equilibrium statistical physics, the study of topological states in discrete stochastic models provides a theoretical approach to predict robust and dynamic function in living systems. The origins of topological classification stem from random matrix theory [7, 8, 9]. These establish the necessary conditions for the existence of models with boundary responses that are robust to generic or random perturbations [10, 11]. In practice, such models are composed of repeated and common motifs (unit cells). This allows for the flexible design of networks that exhibit behavior on a submanifold of the larger system, without requiring fine-tuning or over-parametrization for this specialized behavior on the submanifold.
Such topological invariants have proved useful for analyzing emergent function as they characterize a property of the entire system [10, 12, 13], i.e., whether distinct chemical or mechanical configurations have a global pattern of transition rates between them that supports edge states. Edge states reduce the system response to a lower dimensional space and offer a mechanism for the emergence of global cycles within a large configuration space. Powerfully, this edge response is insensitive to local details, disorder, and noise, and can describe robust biological oscillations or growth cycles. Such theories could shed light on the fundamental question of why biological function is so robust, e.g. during development or when maintaining stable dynamics over long times, even in the presence of stochasticity or changing external conditions and stimuli.
@@
Topological approaches have gained popularity in describing biological systems such as knots in DNA [14] and defects in tissues [15, 16]. Among these approaches, the framework of topological invariants can identify the global state of the system and its associated response, for large classes of different systems [10, 17, 18, 19], see Fig. 1. This framework has led to radical discoveries in other systems, most notably in quantum electrons [10, 12, 17, 20, 21, 22, 23], such as the discovery of dissipationless edge states that enabled precise measurements enabling new standards in metrology [18]. This framework has since been extended to various classical systems, from mechanical lattices [24, 25, 26] to photonics [27, 28, 29, 30], acoustics [31, 32], and active matter [33, 34, 35, 36, 37]. Most of these systems show high spatial order, or even an actual lattice structure in space, which is typically absent in biological systems that are very heterogeneous.
Recently, new mappings of topological tools for non-equilibrium stochastic systems have been demonstrated, and shown to yield localization [35, 34] and currents [33] on the boundaries of an abstract reaction or configuration space. They have been used to model dimensional reduction in the circadian rhythm [33, 40], sensory adaptation and kinetic proofreading [35], and population dynamics [36, 37]. Such models also provide sufficient building blocks for the design of synthetic biological or robotic systems [33].
@@
In this review, we discuss different models that have been proposed in recent years, along with a discussion of attractive properties of these models as well as the formalism used to describe them. In turn, this burgeoning field of study has also prompted the development of new theoretical and computational approaches, as well as the discovery of new physical phenomena. We survey these and discuss the relevant mathematical approaches. We close with a discussion of potential applications to biological and non-biological systems, which herald exciting future directions for the field.
@@
The review is organized as follows. In Section II, we give a brief overview of the development of topological phases initially for quantum matter, and later for classical systems in continuous, real space. In Section III, we describe recent discoveries of topological phases in discrete classical systems, where the dynamics take place in an abstract configuration space. These include stochastic dynamics in biochemical systems, and nonlinear deterministic dynamics in ecological systems. In the remainder of the review, we expand on the novel properties (Section IV), applications (Section V), and theoretical insights and tools (Section VI) that arise in stochastic systems in particular, although much of what we discuss also applies to population dynamics models.
@@ II.1 Topological phases in quantum systems
Topology is concerned with the properties of objects that remain invariant under continuous deformations. While this can sound abstract, the study of topology has made powerful contributions to various physical problems, by distinguishing broad classes of materials and their resulting physical response through analysis of their fundamental properties. Its contribution to the field of quantum electronic systems has been striking, predicting which classes of materials exhibit currents or polarization localized to just the edges of the system [10, 12, 13, 19]. These edge currents are insensitive to local information such as impurities or defects, hence displaying such precise values that new standards for the measurement of fundamental constants have been established [18]. Due to this robust protection from disorder or noise, topological states have also been touted as candidates for fault-tolerant quantum computing [41].
The framework for topological states has been made rigorous thanks to the identification of topological invariants that govern the system. This is due to the celebrated bulk-boundary correspondence, where properties of the bulk (system interior) predict a specific and quantitative response on the boundary [10, 12, 19, 42]. Indeed, materials can be characterized just based on their dimensions and symmetry group [12], through systematic classifications which have been extended to interacting systems [43, 44]. Despite these elegant and fundamental theoretical predictions, the accessibility and detection of topological states can be challenging in practice. Very clean, cold, and ordered samples are often required to observe these delicate quantum properties [10, 13].
@@ II.2 Topological phases in classical systems
Topological approaches in biology have mostly been based on identifying winding structures in biological data, including topological defects in tissues [16, 45, 46, 15, 47] or knots in DNA strands [14], and more abstractly in any high-dimensional dataset through methods such as persistent homology [48, 49, 50, 51, 52, 53]. Other approaches have linked the geometric phase of a system with its dynamics or response, in the context of locomoting or rotating animals [54, 55, 56] or stochastic chemical reactions [57].
The last decade has seen diverse proposals for realizations of new phases of matter characterized by topological invariants in classical platforms. These display edge localization and currents analogous to those in quantum systems, but can be probed in much less extreme conditions. Examples include mechanical lattices [24, 25, 26], photonics [27, 28, 29, 30], acoustics [31, 32], and active matter [33, 34, 35].
@@ III Topological phases in discrete classical systems
Repeated processes or motifs in configuration space, describing a basic reaction such as the conversion of biomolecules from one state to another, or the birth or death of individuals in a population, can be linked together to form an ordered network or lattice that allows for topological analysis. Most importantly, when many such motifs are linked together, it becomes possible to distinguish between the “bulk” of the lattice and its “edges”, the latter arising e.g. from a limiting supply of building blocks or other constraints in the reaction dynamics, that limit the extent of the configuration space. This mapping has been successfully exploited in recent years to allow for the rigorous analysis of topological invariants in stochastic and biological systems, crucially extending models for equilibrium systems to a range of phenomena that is strongly non-equilibrium. They provide a description for the emergence of robust global responses from many basic building blocks (see Fig. 2, top row).
A common example for a building block in the biochemical context is that of futile cycles [38, 39]; see Fig. 2, top left. Futile cycles consist of a series of chemical reactions that bring the molecule back to its starting point, a non-equilibrium process which dissipates energy and thus which requires the input of chemical fuel, typically provided by ATP or GTP in the biological context. Despite seeming wasteful (or futile, as their name suggests), such cycles can be found in protein synthesis, muscular contraction, metabolism, and sensory systems – where their purpose across all these disparate systems remains unclear [39]. Intriguingly, the stochastic topological models that we describe below show that systems with futile cycles as building blocks tend to display topological states [35, 34, 33] . This suggests a fascinating route by which seemingly wasteful non-equilibrium motifs in biology can collectively foster robust global function.
@@ III.1 Stochastic systems
The first demonstration of topological states in stochastic systems was given for one-dimensional (1D) lattices in Ref. 35 and followed-up on shortly after in Ref. 34. A basic depiction of the model and results is shown in Fig. 2(a). The basic lattice has a “ladder” shape, with unit cells containing two sites (top and bottom) that are replicated laterally along the horizontal direction to form a 1D bulk. Considering the transitions (both ways) among internal sites of the unit cell, and among top sites and bottom sites in neighboring cell, a given bulk is defined by six transition rates. The authors considered what occurs at the interface between two different lattices (defined by two different bulks, i.e. two different sets of transition rates) when the two are linked together.
The authors showed that, for any given bulk, one can define a topological index, described further below. Crucially, when the topological indexes of the two bulks that are brought into contact do not coincide, localization (i.e. accumulation of probability at steady state) is observe at the interface between the two.
In summary, these seminal works demonstrated that topological considerations based purely on bulk properties can be used to predict global effect, occurring at the edges of a lattice or at the interfaces between two connected lattices. However, because of the 1D nature of the systems, the topological phenomena that could be observed were limited to static localization.
@@
These restrictions were lifted by us in Ref. 33, which provided the first realization of topological states in two-dimensional (2D) stochastic systems. In 2D systems, besides static polarization, topologically-protected chiral currents along the edges of the system, analogous to those seen in the quantum Hall effect [19], become possible.
@@
Importantly, when (i) internal and external cycles have opposite chirality and (ii) internal transitions are much slower than external ones, probability accumulates at the edges of the system, and a probability current is observed along these edges, with the same chirality as that of the internal cycles. Thus, global oscillations are observed at a system-spanning scale. Further analysis, described below, shows that this effect is topological and thus the edge currents are protected against disorder and deformations of the system boundary.
@@ III.2 Population dynamics
As in the case of biochemical systems, the first demonstration of topological states in population dynamics was given for a 1D lattice in Ref. 36. The authors linked together many rock-paper-scissors cycles into a long chain, see Fig. 2(c). The bulk of the system is thus defined by the three interaction strengths that define the cycle. By construction, in each cycle there is always one ‘isolated’ species that only interacts internally with the other two species in the cycle, and two ‘connected’ species that also take part in the neighboring cycle, and thus interact with four species in total. The authors found that, depending on the skewness of the cycle, determined by the ratio of interaction strengths between the isolated and the two connected species, population would accumulate on either the left or the right end of the chain. Thus, similarly to the 1D biochemical systems, static polarization is observed at steady state. Importantly, polarization was robust to initial conditions and to small disorder in the system parameters, suggesting a topological origin of the transition. This was confirmed by a mapping of the interaction matrix to a Hermitian Hamiltonian that could be analyzed using standard tools of quantum topology, described further below. We note as well that the effect of deviations from ideal rock-paper-scissors interactions in the 1D chain was recently studied in Ref. 69.
@@
In Ref. 37, a 2D Kagome lattice of rock-paper-scissors cycles was studied, see Fig. 2(d). All interactions were chosen to be identical, so that the system is essentially parameter-free. The authors found that, when the system is initialized with additional population in one of the edge sites of the lattice, this additional population propagates along the edges of the lattice, with the same chirality as that of the rock-paper-scissors cycles. As in 2D biochemical systems, this edge current was seen to be topologically protected, and in particular to be robust to deformations of the shape of the boundaries. By mapping the interaction matrix to a Hermitian Hamilonian in a similar way as for the 1D chain, they could show that the edge current is induced by a nontrivial topological index (Chern number).

### v2

@@ Abstract
Topological invariants have proved useful for analyzing emergent function as they characterize a property of the entire system, and are insensitive to local details, disorder, and noise. They support boundary states, which reduce the system response to a lower dimensional space and, in 2D systems, offer a mechanism for the emergence of global cycles within a large phase space. Topological invariants have been heavily studied in quantum electronic systems and have been observed in other classical platforms such as mechanical lattices. However, this framework largely describes equilibrium systems within an ordered crystalline lattice, whereas biological systems are often strongly non-equilibrium with stochastic components. We review recent developments in topological states in discrete stochastic models in 1D and 2D systems, and initial progress in identifying testable signature of topological states in molecular systems and ecology. These models further provide simple principles for targeted dynamics in synthetic systems and in the engineering of reconfigurable materials. Lastly, we describe novel theoretical properties of these systems such as the necessity for non-Hermiticity in permitting edge states, as well as new analytical tools to reveal these properties. The emerging developments shed light on fundamental principles for non-equilibrium systems and topological protection enabling robust biological function.
@@ I Introduction
Understanding how the underlying components of proteins, RNAs, metabolites, and other biologically relevant molecules give rise to biological function would enable precise targeting of interventions to enhance health. However, we still do not have a good theory for how structure leads to function in these complex and noisy systems. As just one example, despite successful sequencing of the human genome, it remains challenging to predict the behavior of resulting proteins and macromolecules due to the large space of possible configurations and reactions causing transitions between them [1, 2]. This large phase space of possible configurations also renders exhaustive searches using other approaches like experiment or numerical simulation unfeasible [3], underscoring the need for simple conceptual methods to provide insight [4, 5, 6]. This review addresses how, drawing on recent theoretical advances in topology, molecular biology, and non-equilibrium statistical physics, the study of topological states in discrete stochastic models provides a theoretical approach to predict robust and dynamic function in living systems.
In this context, topology refers to the study of global features of the system’s eigenvalues and eigenfunctions, particularly in momentum (Fourier) space, that remain unchanged upon continuous deformation. The origins of topological classification stem from random matrix theory [7, 8, 9]. Topological invariants such as winding numbers can be defined that signal topological changes in these mathematical objects, which in turn manifest as phase transitions in the physical system, most famously leading to the emergence of boundary responses that are robust to generic or random perturbations, also known as edge states [10, 11], see Fig. 1. Such topological invariants have proved useful for analyzing emergent function as they characterize a property of the entire system [10, 12, 13].
@@
Traditionally having been used in the quantum condensed matter field, the framework of topological invariants can identify the global state of the system and its associated response, for large classes of different systems [10, 29, 30, 31]. In quantum electronic systems [10, 12, 29, 32, 14, 33, 34], most notably, this framework enabled the discovery of dissipationless edge states that facilitate precise measurements enabling new standards in metrology [30]. This framework has since been extended to various classical systems, from mechanical lattices [15, 16, 17] to photonics [18, 19, 20, 21], acoustics [22, 23], and active matter [24, 26, 25, 27, 28]. Regarding terminology, we must note that the Fourier-space, global topological phenomena described in this review are distinct from real-space, local topological phenomena such as knots in DNA [35] and orientational defects in liquid crystals and biological tissues [36, 37].
Recently, new applications of topological tools to non-equilibrium stochastic systems have been demonstrated, and shown to identify localization [25, 26] and currents [24] on the boundaries of an abstract reaction or configuration space. These systems are often composed of repeated motifs (unit cells) that allow for the flexible design of lattices with edge states, see Fig. 2. The edge states manifest as distinct localization or dynamics on a lower dimensional submanifold of the larger system, without requiring fine-tuning or over-parametrization for this specialized behavior. Powerfully, this edge response is insensitive to local details, disorder, and noise, and can describe robust biological oscillations as global cycles around the boundaries of a large two-dimensional configuration space. They have been used to model dimensional reduction in the circadian rhythm [24, 38], sensory adaptation and kinetic proofreading [25], population dynamics [27, 28], and as building blocks for the design of synthetic biological or robotic systems [24]. These theories could shed light on the fundamental question of why biological function is so robust, e.g. during development or when maintaining stable dynamics over long times, even in the presence of stochasticity or changing external conditions and stimuli.
@@
In this review, we present different topological models that have been proposed in recent years for stochastic systems, along with a discussion of attractive properties of these models as well as the formalism used to describe them. This burgeoning field of study has also prompted the development of new theoretical and computational approaches, as well as the discovery of new physical phenomena. We survey these and discuss the relevant mathematical approaches. We close with a discussion of potential applications to biological and non-biological systems, which herald exciting future directions for the field.
@@
The review is organized as follows. In Section II, we give a brief overview of the development of topological phases initially for quantum matter, and later for classical systems in continuous, physical space. In Section III, we describe recent discoveries of topological phases in discrete classical systems, where the dynamics take place in an abstract configuration space. These include stochastic dynamics in biochemical systems, and nonlinear deterministic dynamics in ecological systems. In the remainder of the review, we expand on the novel properties (Section IV), applications in molecular biology and synthetic systems with potential experimental realizations (Section V), and theoretical insights and tools (Section VI) that arise in stochastic systems in particular, although much of what we discuss also applies to population dynamics models.
@@ II.1 Topological phases in quantum systems
Topology is concerned with the properties of objects that remain invariant under continuous deformations. While this can sound abstract, the study of topology has made powerful contributions to various physical problems, by distinguishing broad classes of materials and their resulting physical response through analysis of their fundamental properties. Its contribution to the field of quantum electronic systems has been striking, predicting which classes of materials exhibit currents or polarization localized to just the edges of the system [10, 12, 13, 31]. These edge currents are insensitive to local information such as impurities or defects. Due to this robust protection from disorder or noise, topological states have also been touted as candidates for fault-tolerant quantum computing [39].
Here, we specifically refer to the algebraic topological features of an object, which can be quantified using topological invariants that undergo discrete changes with the system parameters when one crosses from a topologically trivial phase to a topologically non-trivial phase. Indeed, several invariants have been identified for many different systems that satisfy the Schrödinger equation . Fig. 1 illustrates one example of a topological invariant, in this particular case a winding number.
Because topological models are typically composed of repeating motifs (unit cells) to form lattices, they can be Fourier transformed onto a periodic domain . In the Bloch sphere, the transformed Hamiltonian can wind around the origin as it traverses space, where the number of times it winds around the origin gives the winding number. This winding number is defined as
where is known as the Berry connection and is the eigenvector of for a given band [31]. This is also known as the geometric phase, or the Zak phase [14].
On the left of Fig. 1, we depict systems with a nonzero winding number and hence in a topological phase. Due to the celebrated bulk-boundary correspondence [10, 12, 40] between the lattice properties and its boundaries, this means that the original system (beige line) will host states on its boundary (red dots). Conversely on the right, systems which do not having a winding number are in the trivial phase, where there is no particular distinction between the system bulk and its boundary (beige line).
Another example of a topological invariant is the Chern number,
where bold script denotes a vector and the integral is taken over the two-dimensional Brillouin zone. The Chern number gives the transverse Hall conductance up to a factor of , which is such a precise and robust measurement that it has established a new standard for obtaining fundamental constants of nature [30].
More generally, the number and type of topological invariants that a material can have have been characterized just based on its dimensions and symmetry group [12]. Such systematic classifications were later extended to interacting systems [41, 42]. However, despite these elegant and fundamental theoretical predictions, the accessibility and detection of topological states can be challenging in practice. Very clean, cold, and ordered samples are often required to observe these delicate quantum properties [10, 13].
@@ II.2 Topological phases in classical systems
Topological approaches in biology have mostly been based on identifying local winding structures in biological data, including topological defects in tissues [37, 43, 44, 36, 45] or knots in DNA strands [35], and more abstractly in any high-dimensional dataset through methods such as persistent homology [46, 47, 48, 49, 50, 51]. Other approaches have linked the geometric phase of a system with its dynamics or response, in the context of locomoting or rotating animals [52, 53, 54] or stochastic chemical reactions [55].
The last decade has seen diverse proposals for realizations of new phases of matter characterized by global topological invariants in classical platforms. These display edge localization and currents analogous to those in quantum systems, but can be probed in much less extreme conditions. Examples include mechanical lattices [15, 16, 17], photonics [18, 19, 20, 21], acoustics [22, 23], and active matter [24, 26, 25].
@@ III Topological phases in discrete classical systems
Repeated processes or motifs in configuration space, describing a basic reaction such as the conversion of biomolecules from one state to another, or the birth or death of individuals in a population, can be linked together to form an ordered network or lattice that allows for topological analysis. Most importantly, when many such motifs are linked together, it becomes possible to distinguish between the “bulk” of the lattice and its “edges”, the latter arising e.g. from a limiting supply of building blocks or other constraints in the reaction dynamics, that limit the extent of the configuration space.
For instance, this configuration space could be indexed by the numbers of repeated reactions, or the number of molecular subunits added or proteins transcribed. The repeated motifs would form the bulk, while edges are given by the lowest or highest numbers of the repeating index such as zero reactions or all possible reactions having taken place. Once such a configuration space has been mapped out, we can then apply onto the lattice a rigorous analysis of topological invariants, crucially extending models for equilibrium systems to a range of phenomena that is strongly non-equilibrium. They provide a description for the emergence of robust global responses from many basic building blocks (see Fig. 2).
A common example for a building block in the biochemical context is that of futile cycles [56, 57]; see Fig. 2, top left. Futile cycles consist of a series of chemical reactions that bring the molecule back to its starting point, a non-equilibrium process which dissipates energy and thus requires the input of chemical fuel, typically provided by ATP or GTP in the biological context. Despite seeming wasteful (or futile, as their name suggests), such cycles can be found in protein synthesis, muscular contraction, metabolism, and sensory systems – where their purpose across all these disparate systems remains unclear [57]. Intriguingly, the stochastic topological models that we describe below show that systems with futile cycles as building blocks tend to display topological states [25, 26, 24] . This suggests a fascinating route by which seemingly wasteful non-equilibrium motifs in biology can collectively foster robust global function.
@@ III.1 Stochastic systems
The first demonstration of topological states in stochastic systems was given for one-dimensional (1D) lattices in Ref. 25 and followed-up on shortly after in Ref. 26. A basic depiction of the model and results is shown in Fig. 2(a). The basic lattice has a “ladder” shape, with unit cells containing two sites (top and bottom) that are replicated laterally along the horizontal direction to form a 1D bulk. Considering the transitions (both ways) among internal sites of the unit cell, and among top sites and bottom sites in neighboring cell, a given bulk is defined by six transition rates. The authors considered what occurs at the interface between two different lattices (defined by two different bulks, i.e. two different sets of transition rates) when the two are linked together.
The authors showed that, for any given bulk, one can define a topological index, described in Section III.3. Crucially, when the topological indexes of the two bulks that are brought into contact do not coincide, localization (i.e. accumulation of probability at steady state) is observed at the interface between the two. Intuitively, when the topological invariants do not match, the left bulk and the right bulk have a tendency to create net currents pointing in opposite directions to each other, therefore causing accumulation either at the interface between the two bulks, or at opposite ends of the system.
In summary, these seminal works demonstrated that topological considerations based purely on bulk properties can be used to predict global effects, occurring at the edges of a lattice or at the interfaces between two connected lattices. However, because the net current between unit cells in a finite 1D system is necessarily zero at steady state, the topological phenomena that could be observed were limited to static localization.
@@
These restrictions were lifted by us in Ref. 24, which provided the first realization of topological states in two-dimensional (2D) stochastic systems. In 2D systems, besides static polarization, topologically-protected chiral currents along the edges of the system, analogous to those seen in the quantum Hall effect [31], become possible.
@@
Importantly, when (i) internal and external cycles have opposite chirality and (ii) internal transitions are much slower than external ones, probability accumulates at the edges of the system, and a probability current is observed along these edges, with the same chirality as that of the internal cycles. Thus, global oscillations are observed at a system-spanning scale. Further analysis, described in Section III.3, shows that this effect is topological and thus the edge currents are protected against disorder and deformations of the system boundary.
@@ III.2 Population dynamics
As in the case of biochemical systems, the first demonstration of topological states in population dynamics was given for a 1D lattice in Ref. 27. The authors linked together many rock-paper-scissors cycles into a long chain, see Fig. 2(c). The bulk of the system is thus defined by the three interaction strengths that define the cycle. By construction, in each cycle there is always one ‘isolated’ species that only interacts internally with the other two species in the cycle, and two ‘connected’ species that also take part in the neighboring cycle, and thus interact with four species in total. The authors found that, depending on the skewness of the cycle, determined by the ratio of interaction strengths between the isolated and the two connected species, population would accumulate on either the left or the right end of the chain. Thus, similarly to the 1D biochemical systems, static polarization is observed at steady state. Importantly, polarization was robust to initial conditions and to small disorder in the system parameters, suggesting a topological origin of the transition. This was confirmed by a mapping of the interaction matrix to a Hermitian Hamiltonian that could be analyzed using standard tools of quantum topology, described further below. We note as well that the effect of deviations from ideal rock-paper-scissors interactions in the 1D chain was recently studied in Ref. 69.
@@
In Ref. 28, a 2D Kagome lattice of rock-paper-scissors cycles was studied, see Fig. 2(d). All interactions were chosen to be identical, so that the system is essentially parameter-free. The authors found that, when the system is initialized with additional population in one of the edge sites of the lattice, this additional population propagates along the edges of the lattice, with the same chirality as that of the rock-paper-scissors cycles. As in 2D biochemical systems, this edge current seemed to be topologically protected, as it was found to be robust to deformations or defects in the shape of the boundaries. By mapping the interaction matrix to a Hermitian Hamilonian in a similar way as for the 1D chain, they could show that the edge current is induced by a nontrivial topological index (the Chern number, see Eq. 2).